Circular No. 6602

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The High-Z Supernova Search Team (IAUC 6160) reports the discov-

ery of nineteen supernovae on CCD images taken at the Victor Blanco 4-m telescope at Cerro Tololo by B. Schmidt, L. Germany, N. Suntzeff, R. Schommer, A. Clocchiatti, M. Phillips, and A. Diercks. The supernovae were identified by comparing the images taken in Mar. with those obtained on Feb. 12–13, and confirmed by observing at multiple epochs.

SN  1997 UT  α_{2000}  δ_{2000}  R  z type
1997as Mar. 8 8^{h}24^{m}09.^{s}46 − 0^{o}48′08′′.2 22.1 0.47? Ia?
1997at Mar. 8 10 43 37.63 − 0 49 06.5 21.6
1997au Mar. 8 11 50 04.24 +10 45 22.4 23
1997av Mar. 8 10 23 11.24 + 4 13 44.5 21.4
1997aw Mar. 8 10 23 28.21 + 4 07 11.1 22.7 0.6? ?
1997ax Mar. 8 11 18 02.05 + 0 02 10.5 21.8 ? ?
1997ay Mar. 8 11 26 10.34 − 0 36 26.9 23 0.62 ?
1997az Mar. 8 11 23 05.57 − 0 15 55.6 19 0.04 II
1997ba Mar. 8 12 27 38.34 + 0 13 31.0 22.5
1997bb Mar. 8 12 29 00.72 + 0 09 11.3 22.1
1997bc Mar. 8 11 59 08.27 − 0 33 20.6 22.1 0.26? II
1997bd Mar. 8 12 40 28.75 − 0 05 41.8 23.1
1997be Mar. 8 12 40 40.78 − 0 23 21.8 23.1
1997bf Mar. 8 13 41 48.35 + 0 06 12.0 20.4 0.15 ?
1997bg Mar. 8 12 40 39.52 − 0 24 02.7 23.5 0.19 II
1997bh Mar. 8 13 44 35.97 − 0 19 49.3 22.3 0.40 Ia
1997bi Mar. 9 10 11 28.88 −13 12 55.4 23.3
1997bj Mar. 9 10 42 26.65 + 0 01 35.4 21.8 0.35 Ia
1997bk Mar. 9 10 42 02.11 + 0 01 48.2 22.4 0.28? II

Each supernova is within 3′′ of its host galaxy’s nucleus, except for SNe 1997aw, 1997be, and 1997bh, for which no host galaxies are clearly identified. For the supernovae with indicated redshift and type, spectra were obtained with the European Southern Observatory 3.6-m telescope by J. Spyromilio on Mar. 12; and with the Multiple-Mirror Telescope on Mar. 11–12 by R. Kirshner and S. Jha. Spectral reductions were performed by B. Leibundgut and P. Challis. In addition, a spectrum of SN 1997bh was obtained with the Keck telescope by S. Perlmutter, R. Knop, and I. Hook. Finder charts may be requested by sending electronic mail to

pgarnavich@cfa.harvard.edu.

1997 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green